A field is in the shape of a trapezium whose parallel sides are 25m and 10m the non-parallel Sides are 14m and 13m,Find the area of field

Grade:9

1 Answers

Rajat
5 years ago
Let ABCD be the trapezium. And AB= 10 CD=25 BC=13 AD=14. 
Let AE and BF be perpendicular on CD, so AE=BF
Triangle ADE, BCF are right angled.
AB=EF=10. Let CF = X so, DE = 15-X 
By Pythagoras Theorem 142+(X-15)2= 132+(X)2
so X comes out to be 252/30= 8.4 
So, AE= BF = √(132-8.42) = 9.92
So area of trapezium = 1/2(10+25)(9.92)= 173.6